OpenStudy (anonymous):

What is common to the variation in the range and the height of a projectile? a - horizontal velocity b- time of flight vertical velocity, horizontal acceleration c - vertical velocity d - horizontal acceleration

OpenStudy (anonymous):

I am considering you are talking about the maximum height of the projectile then Range = (v^2*sin2θ)/g & Hmax=(v^2*sin^2θ)/2g We see that Vsinθ that is vertical velocity is common in both